OpenStudy (et365):

If it has the vertex (1,-4) and passes through the point (2,3) what is this in quadratic form?

OpenStudy (anonymous):

Formula is f(x)=a(x-h)^2+k

OpenStudy (et365):

That is in vertex form, I understand that, but I'm not too sure what it is in quadratic form

OpenStudy (anonymous):

Points are (x,y) =(2,3) Vertex is (h,k) = (1,-4) 3=a(2-(1))^2+(-4) Do the parentheses first which is 3=a(1)^2 +(-4) Then do the exponent 3=a(1)+(-4) Add 4 to both sides 7=a Then use the formula and plug it in So the answer should be F(x)=7(x-1)-4

OpenStudy (campbell_st):

well when you find a, but substituting the point you can then distribute to get the equation

OpenStudy (anonymous):

This is the standard form for quadratics

OpenStudy (campbell_st):

and it should be \[f(x) =7(x -1)^2 - 4 \]

OpenStudy (anonymous):

I forgot to put the square in sorry

OpenStudy (et365):

Hmm so any way to convert that to non-standard?

OpenStudy (campbell_st):

so you want it as \[f(x)=ax^2 + bx + c\] is that correct...?

OpenStudy (et365):

Yes

OpenStudy (campbell_st):

ok... well its \[f(x) = 7(x^2 -2x + 1) - 4\] just distribute the 7 and collect like terms

OpenStudy (et365):

Thanks, I was unsure of which form of the equation to sub a in.